Three particles are fixed on an x axis. Particle 1 of charge q1 is at x=-a and particle 2 of charge q2 is at x=+a. If their net electrostatic force on particle 3 of charge Q is to be zero, what must be hte ratio q1/q2when particle 3 is at (a) x=+0.513a and (b) x=+1.22a?
Answer #1


(a) q1 / 1.513 ^ 2 = q2 / (1 - 0.513)^2 ==> q1/q2 = (1.513 / 0.487)^2 = 9.652 (Answer)

(b) q1 / 2.22 ^ 2 = -q2 / 0.22^2 ==> q1/q2 = -(2.22 / 0.22)^2 = -101.826 (Answer)